# Knapsack Optimization-based Schema Linking for LLM-based Text-to-SQL Generation

### v1

@@ Abstract
Generating SQLs from user queries is a long-standing challenge, where the accuracy of initial schema linking significantly impacts subsequent SQL generation performance. However, current schema linking models still struggle with missing relevant schema elements or an excess of redundant ones. A crucial reason for this is that commonly used metrics, recall and precision, fail to capture relevant element missing and thus cannot reflect actual schema linking performance. Motivated by this, we propose an enhanced schema linking metric by introducing a restricted missing indicator. Accordingly, we introduce Knapsack optimization-based Schema Linking Agent (KaSLA), a plug-in schema linking agent designed to prevent the missing of relevant schema elements while minimizing the inclusion of redundant ones. KaSLA employs a hierarchical linking strategy that first identifies the optimal table linking and subsequently links columns within the selected table to reduce linking candidate space. In each linking process, it utilize a knapsack optimization approach to link potentially relevant elements while accounting for a limited tolerance of potential redundant ones. With this optimization, KaSLA-1.6B achieves superior schema linking results compared to large-scale LLMs, including deepseek-v3 with state-of-the-art (SOTA) schema linking method. Extensive experiments on Spider and BIRD benchmarks verify that KaSLA can significantly improve the SQL generation performance of SOTA text-to-SQL models by substituting their schema linking processes.
Knapsack Optimization-based Schema Linking for
LLM-based Text-to-SQL Generation
Zheng Yuan1, Hao Chen2, Zijin Hong1, Qinggang Zhang1 Feiran Huang3, Xiao Huang1 1The Hong Kong Polytechnic University 2City University of Macau, 3Jinan University {yzheng.yuan, zijin.hong, qinggangg.zhang}@connect.polyu.hk sundaychenhao@gmail.com; huangfr@jnu.edu.cn; xiao.huang@polyu.edu.hk
1 Introduction
@@
Figure 1: In text-to-SQL tasks, commonly used schema linking metrics such as Recall and Precision fail to accurately capture the true performance of schema linking when relevant schema elements are missing.
Figure 2: Overall framework of proposed knapsack optimization-based schema linking agent (KaSLA) with column linking as an example. Traditional schema linking models often struggle with both missing relevant elements and the inclusion of redundant ones. KaSLA addresses this challenge by maximizing the total estimated relevance while maintaining total redundancy within a specified tolerance using knapsack optimization. Serving as a general model, KaSLA enhances the SQL generation of text-to-SQL models by replacing their existing schema linking processes.
With the advent of large language models (LLMs) (Achiam et al., 2023; Dubey et al., 2024), LLM-based text-to-SQL is emerging as the next-generation interface for database users (Hong et al., 2024b). Typically, text-to-SQL frameworks employ a two-step process: first linking database schema elements (tables and columns) to user queries, then generating the corresponding SQL statements. However, the accurate SQL generation requires the linking to the correct schema elements, meaning that errors in early stages are inevitably propagated to later SQL generation. For instance, inaccurate schema linking—such as overlooking relevant schema elements or linking to irrelevant ones—will lead to incorrect SQL statements.
Recent state-of-the-art text-to-SQL models tend to focus primarily on final SQL generation while employing relatively simplistic schema linking strategies or neglecting it. DTS-SQL (Pourreza and Rafiei, 2024) finetunes a local LLM to select relevant tables based on user queries. Solid-SQL (Liu et al., 2025) present a data augmentation method for the finetuning of Schema linking method. Chess (Talaei et al., 2024) and CodeS (Li et al., 2024b) use semantic-based retrieval models to enhance the schema linking process by retrieving relevant information from databases. RSL-SQL (Cao et al., 2024) introduces a bidirectional approach to simplify the full schema. E-SQL (Qu et al., 2024) and (Maamari et al., 2024) assert that schema linking remains challenging for current LLMs and thus directly use the full schema as input for SQL generation.
Even though schema linking is crucial, current widely used metrics, such as Precision and Recall (Li et al., 2023a, 2024b; Qu et al., 2024), cannot adequately evaluate instances when relevant elements are missing. As demonstrated in Figure 1, missing a single relevant table or column can significantly decrease the accuracy of SQL generation, yet still result in high Recall and Precision scores. Because Recall does not significantly penalize the model as long as the majority of the relevant elements are identified, allowing it to achieve an artificially high score, even when crucial elements are missing. Similarly, regardless of the severity of the missing, as long as the number of correctly linked elements is proportionally high compared to the total predicted relevant elements, the model can achieve an unreasonably high Precision score. This discrepancy distorts the assessment of a schema linking model’s effectiveness, potentially hindering progress in text-to-SQL development.
@@
To address this, we firstly design a tailored schema linking metrics that accurately assess linking performance with a restricted element missing indicator. Additionally, we propose the Knapsack optimization-based Schema Linking Agent (KaSLA), designed to link the most relevant and minimally redundant schema elements according to the given queries. KaSLA introduces a binary-probabilistic score function to predict a robust relevance score of each element. Additionally, KaSLA identifies redundancy of each element and the upper redundancy tolerance for each query based on its most similar queries in the training dataset, effectively minimizing the inclusion of redundant elements. Finally, KaSLA links tables and columns with an hierarchical strategy to reduce the column linking candidate space for efficiency and employs a 0-1 knapsack optimization approach to select the most relevant element set under the upper redundancy tolerance to reduce the redundant linking. Extensive experiments demonstrate that KaSLA can improve SQL generation accuracy by reducing missing and redundant information in the schema linking.
@@
  • We introduce an enhanced schema linking metric, which can accurately capture the missing of relevant elements and reflect the actual linking performance aligning with the SQL generation performance.
  • We propose the Knapsack optimization-based Schema Linking Agent (KaSLA) to ensure the linking of relevant schema elements and minimizing the inclusion of redundant ones.
  • KaSLA can serve as a general agent to enhance LLM-based text-to-SQL framework by provide optimal linking results. Extensive experiments demonstrate improved performance in the SOTA text-to-SQL models when using KaSLA for schema linking.
2 Preliminaries
In this section, we outline the definitions for the schema linking and 0-1 Knapsack Optimization.
Schema Linking Task.
Given a natural language query $q$ and a corresponding database schema ${\mathcal{S}}_{q}$ containing schema elements, tables ${\mathcal{T}}=\{t_{1},\cdots,t_{|{\mathcal{T}}|}\}$ and columns ${\mathcal{C}}=\{c_{1},\cdots,c_{|{\mathcal{C}}|}\}$ in each table. Schema linking is a common process used in text-to-SQL to predict the relevant schema entities with $q$ from ${\mathcal{S}}_{q}$:
@@ 2 Preliminaries
where $\widehat{{\bm{S}}}_{q}$ is the subset of ${\mathcal{S}}_{q}$ that includes schema element $s\in{\mathcal{S}}_{q}$ that relevant to $q$. Each $s$ could represent a specific table or column. Additionally, we denote ${\bm{S}}_{q}^{*}$ as the ground truth linking results.
0-1 Knapsack Optimization.
The 0-1 knapsack problem Fréville (2004) is a classic combinatorial optimization challenge. It involves selecting items from a finite item set, each with a value ${\mathcal{V}}$ and weight ${\mathcal{W}}$, to place into a knapsack with a maximum weight capacity ${\mathcal{C}}$. Each item can be selected only once. Formally, the objective of 0-1 knapsack optimization is to select items such that their total weight remains below ${\mathcal{C}}$ while maximizing total value of these selected items:
@@
3 Enhanced Schema Linking Metrics
In this section, we introduce enhanced schema linking metrics considering the missing of relevant elements during evaluation, thereby accurately assess the true performance of schema linking models.
For a query $q$ with predicted schema linking result $\widehat{{\bm{S}}}_{q}$ and ground truth linking result ${\bm{S}}^{*}_{q}$, Recall and Precision are two commonly used metreics (Li et al., 2023a, 2024b; Qu et al., 2024). Recall, ${\bm{R}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|{\bm{S}}^{*}_{q}|}$, measures the proportion of relevant elements correctly identified, specifically how many of the actual relevant elements in ${\bm{S}}^{*}_{q}$ are found in $\widehat{{\bm{S}}}_{q}$. Conversely, Precision, ${\bm{P}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|\widehat{{\bm{S}}}_{q}|}$, assesses the accuracy of predictions by determining the proportion of elements predicted as relevant in $\widehat{{\bm{S}}}_{q}$ that are indeed correct.
As illustrated in Figure 1, these metrics can yield high scores even in the missing of relevant elements. This underscores that they do not effectively address the issue of missing relevant elements, which can undermine the true accuracy of schema linking models.
@@ 3 Enhanced Schema Linking Metrics
It takes a value of 1 when the predicted result $\widehat{{\bm{S}}}_{q}$ contain all the elements in the ground truth ${\bm{S}}^{*}_{q}$, meaning no relevant elements are missing. Conversely, it is set to 0 when there are any missing. Building on this missing indicator, we define enhanced Recall and Precision as follows:
@@
Such metrics strengthen the penalty for missing relevant elements, as schema linking results with missing are inherently incorrect for SQL generation. For simplicity, we use the harmonic mean of enhanced Recall and Precision, referred to as the enhanced F1 score, as the overall measure for evaluating schema linking results:
@@
This approach provides a comprehensive evaluation of both recall and precision while incorporating the critical aspect of missing elements.
4 Knapsack Optimization-based Schema Linking Agent (KaSLA)
In this section, we introduce the proposed KaSLA framework, which focuses on ensuring the linking of relevant schema elements to avoid missing and minimizing the inclusion of redundant ones. As illustrated in Figure 2, We begin by detailing a binary-probabilistic score function used for assessing the relevance of schema elements and an estimation function for the redundancy tolerance of each schema linking process. We then outline the knapsack optimization approach to link potential relevant element while reducing the potential redundant ones under the tolerance constraint. Finally, we describe the hierarchical schema linking process, demonstrating its efficiency for reducing the column linking candidate.
4.1 Relevance Estimation

### v2

@@ Abstract
Generating SQLs from user queries is a long-standing challenge, where the accuracy of initial schema linking significantly impacts subsequent SQL generation performance. However, current schema linking models still struggle with missing relevant schema elements or an excess of redundant ones. A crucial reason for this is that commonly used metrics, recall and precision, fail to capture relevant element missing and thus cannot reflect actual schema linking performance. Motivated by this, we propose enhanced schema linking metrics by introducing a restricted missing indicator. Accordingly, we introduce Knapsack optimization-based Schema Linking Approach (KaSLA), a plug-in schema linking method designed to prevent the missing of relevant schema elements while minimizing the inclusion of redundant ones. KaSLA employs a hierarchical linking strategy that first identifies the optimal table linking and subsequently links columns within the selected table to reduce linking candidate space. In each linking process, it utilizes a knapsack optimization approach to link potentially relevant elements while accounting for a limited tolerance of potentially redundant ones. With this optimization, KaSLA-1.6B achieves superior schema linking results compared to large-scale LLMs, including deepseek-v3 with the state-of-the-art (SOTA) schema linking method. Extensive experiments on Spider and BIRD benchmarks verify that KaSLA can significantly improve the SQL generation performance of SOTA Text2SQL models by substituting their schema linking processes.
Index Terms:
text-to-SQL, database, large language models, natural language understanding
I Introduction
With the advent of large language models (LLMs) [1, 2], LLM-based text-to-SQL is emerging as the next-generation database interface for unskilled users [3]. Typically, text-to-SQL frameworks employ a two-stage process: first linking user queries to database schema elements (tables and columns), then generating the corresponding SQL statements. Accurate SQL generation relies on correctly identifying these schema elements, since errors in this initial linking phase inevitably propagate to subsequent SQL construction. Inaccurate schema linking, such as overlooking relevant schema elements or linking to irrelevant ones, consistently produces erroneous SQL. This challenge is shown in Figure 1. When using CodeS-15B, a pre-trained LLM for SQL generation, on the BIRD-dev benchmark, there is a significant 14.91% performance gap between state-of-the-art schema linking methods and ground-truth linking results. Consequently, enhancing schema linking accuracy represents a critical research frontier with significant potential to advance text-to-SQL capabilities.
Figure 1: Performance comparison of different schema linking methods on the BIRD-dev dataset, using CodeS-15B as the SQL generation model.
@@ I Introduction
Figure 2: Commonly used metrics, like Recall and Precision, fail to accurately reflect actual schema linking performance when relevant elements are missing.
Recent state-of-the-art text-to-SQL models primarily focus on the final SQL generation stage, often relying on basic or even neglecting schema linking strategies. Existing schema linking methods can generally be categorized into three types. First, some frameworks directly use the full database schema as input without a specialized schema linking mechanism, instead depending on the reasoning capabilities of LLMs. For example, DAIL-SQL [4], E-SQL [5], and Distillery-SQL [6] enhance LLMs’ ability to process full schema through effective prompting. Second, recall-based schema linking methods have been developed. For instance, ResdSQL [7] and CodeS [8] train encoder-based language models to rank and predict relevant schema elements. They employ decoupled linking strategies to separately handle tables and columns, enabling more effective use of semantic information. Third, generative schema linking methods typically leverage LLMs as schema linking models. DTS-SQL [9] finetunes a local LLM to select relevant tables given user queries, while Solid-SQL [10] proposes data augmentation to improve schema linking performance. Other approaches, such as TA-SQL [5], use in-context learning to guide LLMs in generating dummy SQL that abstracts linked schema elements. Multi-agent models like CHESS [11] design dedicated schema selector agents that prompt LLMs to choose relevant tables and columns. Furthermore, RSL-SQL [12] adopts a bidirectional approach for schema simplification, aiming to further refine schema linking.
Despite the critical role of schema linking, traditional metrics like Recall and Precision [7, 8, 5] fail to assess missing relevant elements adequately. A single missing element can render SQL generation impossible, yet current metrics may still yield high scores that misrepresent true model performance. As shown in Figure 2, the schema linking result $\widehat{{\bm{S}}}_{q}$ misses one key column, organization, which directly causes the failure to generate correct SQL due to incomplete schema linking results. However, traditional Recall and Precision metrics still give $\widehat{{\bm{S}}}_{q}$ falsely high scores. This happens because Recall does not severely penalize the model if most relevant elements are found, allowing a high score despite missing a few relevant elements. Likewise, Precision can be unreasonably high as long as the number of correctly linked elements is large compared to the total predicted ones, regardless of how many are missed. This discrepancy skews schema linking evaluation and may hinder progress in text-to-SQL research.
To address this issue, we propose tailored schema linking metrics, Recall+ and Precision+, which incorporate a restricted missing-element indicator to more accurately assess linking performance. These metrics impose penalties when relevant elements are missing, since the absence of any relevant element in the schema linking results prevents the SQL generation model from obtaining the complete information needed to generate the correct SQL. Based on the proposed evaluation metrics, we find that existing schema linking methods face a significant issue—the Missing & Redundancy Seesaw Problem: There exists a seesaw phenomenon between missing and redundant elements in current schema linking methods: Generative linking strategies tend to have fewer redundant elements but more missing ones, while recall-based models usually have fewer missing elements but more redundancy. It is difficult for current methods to achieve both low missing and low redundancy at the same time, creating a trade-off situation similar to a seesaw. Developing schema linking approaches that can effectively reduce both missing and redundant schema elements remains a significant challenge.
@@
To address this seesaw problem, we propose the Knapsack Optimization-based Schema Linking Aapproach (KaSLA), aims to link the relevant schema elements for each query while minimizing missing and redundant cases simultaneously. KaSLA introduces a binary-probabilistic score function to predict a robust relevance score for each element. In addition, KaSLA identifies redundancy of each element and determines an upper redundancy tolerance for each query by referencing its most similar queries in a pre-constructed pool of query and linking result pairs. Finally, KaSLA performs table and column linking hierarchically; it links columns based on intermediate table linking results, which reduces the candidate search space. During each linking step, KaSLA uses a 0-1 knapsack optimization approach to select the most relevant set of elements under the defined upper redundancy tolerance, aiming to avoid missing elements and minimize redundancy, all within an efficient dynamic programming framework. Extensive experiments show that KaSLA improves a text-to-SQL framework’s SQL generation accuracy by simultaneously reducing missing and redundant information in schema linking.
@@
  • We introduce enhanced schema linking metrics, Recall+ and Precision+, which can accurately capture the missing of relevant elements and reflect the actual linking performance aligning with the SQL generation.
  • We propose the Knapsack optimization-based Schema Linking Aapproach (KaSLA) to link the relevant schema elements for each query while minimizing missing and redundant elements simultaneously.
  • KaSLA can serve as a general plug-in method to enhance LLM-based text-to-SQL framework by provide optimal linking results. Extensive experiments demonstrate improved performance in the SOTA text-to-SQL models when using KaSLA for schema linking.
II Preliminary Study
In this section, we outline the definitions for the schema linking task, SQL generation task and 0-1 knapsack optimization. Additionally, we introduce the traditional schema linking metrics and then discuss their limitations.
II-A Definitions
II-A1 Schema Linking Task
Given a natural language query $q$ and a corresponding database schema ${\mathcal{S}}_{q}$ containing schema elements, tables ${\mathcal{T}}=\{t_{1},\cdots,t_{|{\mathcal{T}}|}\}$ and columns ${\mathcal{C}}=\{c_{1},\cdots,c_{|{\mathcal{C}}|}\}$ in each table. Schema linking is a common process used in text-to-SQL to predict the relevant schema elements (tables and columns) with $q$ from ${\mathcal{S}}_{q}$:
@@ II-A1 Schema Linking Task
where $\widehat{{\bm{S}}}_{q}$ is a subset of ${\mathcal{S}}_{q}$ that includes schema element $s\in{\mathcal{S}}_{q}$ that relevant to $q$ and will be used for subsequent SQL generation. Each $s$ could represent a specific table or column. Additionally, we denote ${\bm{S}}_{q}^{*}$ as the ground truth linking results.
II-A2 SQL Generation Task
With the natural language query $q$ and the schema linking results $\widehat{{\bm{S}}}_{q}$, the function of SQL generation task is to generate an SQL statement that is able to retrieve the pertinent information of $q$ from the database:
$f_{\text{sql}}:(q,\widehat{{\bm{S}}}_{q})\mapsto{\text{S}\widehat{\text{Q}}\text{L}}_{q},$ (2)
where $\text{S}\widehat{\text{Q}}\text{L}_{q}$ represents the predicted SQL statement.
II-A3 0-1 Knapsack Optimization
The 0-1 knapsack problem [13] is a classic combinatorial optimization challenge. It involves selecting items from a finite item set, each with a value ${\mathcal{V}}$ and weight ${\mathcal{W}}$, to place into a knapsack with a maximum weight capacity ${\mathcal{C}}$. Each item can be selected only once. Formally, the objective of 0-1 knapsack optimization is to select items such that their total weight remains below ${\mathcal{C}}$ while maximizing total value of these selected items:
@@ II-A2 SQL Generation Task
II-B Traditional Schema Linking Metrics
II-B1 AUC
AUC is utilized to evaluate schema linking by treating it as a binary classification task, relevant elements are labeled as 1 and irrelevant ones as 0 [5]. However, as depicted in Figure 3, the significant imbalance between relevant and irrelevant elements will hinder the AUC from accurately reflecting the linking performance. A model can cheat its way to a high AUC by classifying many irrelevant elements, an easier task than accurately identifying the few relevant ones.
II-B2 Recall and Precision
Recall and Precision are commonly used in recent text-to-SQL research to evaluate schema linking performance [7, 8, 5]. For a query $q$ with the predicted schema linking result $\widehat{{\bm{S}}}_{q}$, their formula is shown below:
${\bm{R}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|{\bm{S}}^{*}_{q}|},{\bm{P}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|\widehat{{\bm{S}}}_{q}|},$ (4)
where ${\bm{S}}^{*}_{q}$ is the ground truth linking result. Recall ${\bm{R}}(q)$ measures the proportion of relevant elements correctly identified, specifically how many of the actual relevant elements in ${\bm{S}}^{*}_{q}$ are found in $\widehat{{\bm{S}}}_{q}$. Conversely, Precision ${\bm{P}}(q)$ assesses the accuracy of predictions by determining the proportion of elements predicted as relevant in $\widehat{{\bm{S}}}_{q}$ that are indeed correct.
As shown in Figure 2, the predicted schema linking result $\widehat{{\bm{S}}}_{q}$ fails to include an essential column, “organization”. Without this information, the SQL generation model cannot generate the correct SQL statement. Nonetheless, Recall and Precision still report high scores, as they mainly reflect the number of correctly identified elements but overlook the impact of missing relevant ones. This demonstrates that these metrics can overestimate the effectiveness of schema linking, especially in this scenario where even a single omission is detrimental.
Figure 3: The imbalanced proportion of relevant schema elements with the irrelevant ones in the commonly used text-to-SQL datasets.
Figure 4: Missing & redundancy seesaw problem in schema linking. We evaluate the column linking performance of schema linking baselines on BIRD-dev.
III Enhanced Schema Linking Metrics
In this section, we introduce enhanced schema linking metrics considering the absence of relevant elements during evaluation, thereby accurately assessing the performance of schema linking models. Addtionaly, we discuss the missing & redundancy seesaw problem in current schema linking methods.
III-A Recall+ and Precision+:
@@
It takes a value of 1 when the predicted result $\widehat{{\bm{S}}}_{q}$ contain all the elements in the ground truth ${\bm{S}}^{*}_{q}$, meaning no relevant elements are missing. Conversely, it is set to 0 when there are any missing. Building on this missing indicator, we define Recall+ and Precision+ as follows:
@@
Such metrics strengthen the penalty for missing any relevant element, as schema linking results with missing are inherently incorrect for SQL generation.
Figure 5: Overall framework of proposed knapsack optimization-based schema linking approach (KaSLA) and and how it operates within a text-to-SQL system. Traditional schema linking models often struggle with both missing relevant elements and the inclusion of redundant ones. KaSLA addresses this challenge by maximizing the total estimated relevance while maintaining total redundancy within a specified tolerance using knapsack optimization. Serving as a general plug-in method, KaSLA enhances the SQL generation of text-to-SQL models by replacing their existing schema linking processes.
Since Recall+ evaluates whether there are missing elements in the linking results and Precision+ measures the proportion of redundant elements, we use their harmonic mean, called F1+, as the overall metric. The formula is as follows:
@@
F1+ combines the strengths of Recall+ and Precision+ and provides a comprehensive evaluation of schema linking.
Taking the same example in Figure 2, our enhanced metrics Recall+ and Precision+ offer a more reliable evaluation. When the predicted result lacks any key element, such as the “organization” column, these metrics assign a score of zero, signaling a definite failure in schema linking for SQL generation. By strictly penalizing missing crucial information, Recall+ and Precision+ ensure models are only rewarded for comprehensive and complete predictions, thus aligning the metric outcomes closely with actual SQL generation success.
III-B The Missing & Redundancy Seesaw Problem
After eliminating the misleading effects caused by missing elements, Recall+ and Precision+ can be used to accurately measure a schema linking model’s ability to link relevant elements and avoid redundant elements, respectively. We applied several baselines on BIRD-dev and evaluated their column linking preferences. As shown in Figure 4, we observed a significant seesaw phenomenon between missing and redundancy in these schema linking methods.
For example, recall-based models like CodeS-SL rarely miss relevant columns, resulting in high Recall+, but it includes too many redundant columns, leading to very low Precision+. In contrast, generative methods such as Pure-SL, RSL-SL and TA-SL tend to have more missing columns, resulting in lower Recall+, but their linking results contain fewer redundant columns, yielding relatively higher Precision+ compared to the recall-based model, CodeS-SL.
This prominent missing & redundancy seesaw problem inspires a key question in schema linking: How can we minimize both element missing and element redundancy while ensuring that relevant elements are properly linked? To address this issue, we propose the Knapsack Optimization-based Schema Linking Approach (KaSLA), which uses a knapsack optimization framework to improve the schema linking process. The details are provided in Section IV.
IV Knapsack Optimization-based Schema Linking Approach (KaSLA)
In this section, we present KaSLA, a novel schema linking framework for text-to-SQL tasks that explicitly addresses both missing and redundant element challenges through a knapsack optimization approach. We first provide an overview of the KaSLA framework in Section IV-A, outlining its key components, optimization objective and hierarchical structure. Next, we introduce the core components in Section IV-B, including the hybrid relevance estimation mechanism, redundancy assessment strategy, and the formulation of redundancy tolerance for each query. We then describe how schema linking works as a knapsack optimization process and explain the hierarchical linking strategy for efficient and precise schema element selection in Section IV-C. Each subsection details a critical aspect of KaSLA, jointly enabling accurate and efficient schema linking for robust text-to-SQL generation.
IV-A Overall Framework
As illustrated in Figure 5, KaSLA is designed to systematically address the challenges of both missing and redundant schema elements in text-to-SQL systems by using a knapsack optimization approach. The overall process begins with two key inputs: a natural language query and the full database schema. As visible in the left portion of Figure 5, KaSLA follows a hierarchical linking strategy, where it first performs table linking to identify tables relevant to the query, and then proceeds to link columns within each selected table. This approach effectively reduces the candidate space and enhances efficiency when working with complex database schema.
At the core of KaSLA’s schema linking is a hybrid scoring module that combines a binary scoring function (yielding hard 0/1 relevance decisions) and a probabilistic scoring function (providing soft relevance probabilities) for each schema element. As shown in the right side of Figure 5, these scores are integrated to assign a final relevance score $r$ to every candidate element. In parallel, each element receives a redundancy score $w$, and for each query, a redundancy tolerance $u$ is estimated by analyzing similar queries with their linking results from a pre-constructed query-linking pool, as depicted in the figure. This step ensures the model adapts its tolerance to different query types and schema complexities.
The selection of relevant schema elements is then formulated as a 0-1 knapsack problem: KaSLA seeks to maximize the total relevance scores of linked elements while keeping the sum of their redundancy scores within the estimated tolerance $u$. The optimization objective, shown in red in Figure 5, is solved efficiently using a dynamic programming algorithm. The knapsack selection process ensures that the final set of linked schema elements contains all necessary information for SQL generation while minimizing unnecessary redundancy.
After schema linking, the selected elements are passed to the downstream SQL generation model, which constructs the final SQL statement in accordance with the linking result. As shown on bottom left of Figure 5, this process allows KaSLA to serve as a plug-in method for existing text-to-SQL frameworks, helping them avoid common pitfalls of missing relevant elements or introducing redundant ones. By leveraging hierarchical optimization, hybrid scoring and redundancy-aware selection, KaSLA robustly improves schema linking across diverse and large-scale text-to-SQL applications.
IV-B Key Factor Estimation
In this section, we present a detailed introduction to the estimation functions of key factor in KaSLA: element relevance, element redundancy and upper redundancy tolerance.
IV-B1 Relevance Estimation